OpenStudy (anonymous):

Which of the following equations represent the line that has slope 7 and goes through the point (2, 9)? y + 9 = 7(x + 2) y - 2 = 7(x - 9) y - 2 = -7(x - 9) y - 9 = 7(x - 2)

jagr2713 (jagr2713):

do you know the formula for point slope form?

jagr2713 (jagr2713):

y-y1=m(x-x1)

jagr2713 (jagr2713):

m=7 (2,9) x y

jagr2713 (jagr2713):

you there @juciystar234

OpenStudy (anonymous):

y - 2 = -7(x - 9 ) y + 9 = 7(x + 2)